# Arithmetic dynamics and Generalized Fermat’s conjecture

### v3

@@ Abstract.
I would like to thank Prof. Kawaguchi for his helpful comments.
@@ 1. Introduction
Let $K$ be an arithmetic function field (i.e., $K$ is finitely generated over $\mathbb{Q}$}) and $S=((\Omega,\mathcal{A},\nu),\{|\raisebox{1.72218pt}{.}|_{\omega}\}_{\omega\in\Omega})$ be a proper adelic structure of $K$ such that $S$ has Northcott’s property and $\nu(\mathcal{A})\not\subseteq\{0,1\}$ (for the definition of “adelic structure” and its related terminology, see [1]). Let $X$ be a geometrically integral projective scheme over $K$ and $L$ be an ample line bundle on $X$. Fix a positive integer $N_{0}$. Let $F=\{f_{N}\}_{N=N_{0}}^{\infty}$ be a sequence of endomorphisms of $X$ with the following properties:
@@
The sequence $F$ is called a system of endomorphisms compatible with $L$. We say $F$ is multiplicative (resp. additive) if $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{NN^{\prime}}$ (resp. $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{N+N^{\prime}}$) for all $N,N^{\prime}\geqslant N_{0}$. For example, the cases of Example 3.2, Example 3.3, Example 3.4 and Example 3.5 are multiplicative, and the case of Example 3.6 is additive. We can assign the height function $h_{F}$ on $X(K)$ to $F$ such that
@@
    If $F$ is additive and there is $N_{1}\geqslant N_{0}$ such that $Y_{N_{1}}(K)$ is finite, then there exists $N_{2}\geqslant N_{0}$ such that $Y_{N}$ has Fermat’s property for any $N\geqslant N_{2}$. Namely, Conjecture 1.1 holds for an additive system of endomorphisms compatible with $L$.
@@
    $\lim_{m\to\infty}\frac{\#\{N_{0}\leqslant N\leqslant m\mid\text{$F_{N}$ has Fermat's property over $K$}\}}{m}=1.$
2. Endomorphism compatible with an ample line bundle
Let $f:X\to X$ be an endomorphism over $K$. Let $L$ be an ample line bundle on $X$. We say $f$ is compatible with $L$ if there is an isomorphism $\alpha:f^{*}(L)\simeq L^{\otimes d(f)}$ for some positive integer $d(f)$.
@@ 2. Endomorphism compatible with an ample line bundle
    Let $g$ be another endomorphism of $X$ such that $g$ is compatible with $L$ and $d(g)\geqslant 2$. If $f\circ g=g\circ f$, then $h_{f}=h_{g}$. Moreover, $f\circ g$ is compatible with $L$, and $d(f\circ g)=d(f)d(g)$.
@@
(2) For the existence of $\varphi_{f,\alpha}$, we refer to [1, Proposition 2.5.11 (4)]. The second assertion follows from [3, Proposition 9.3.7].
(3) Let $\beta$ be another isomorphism $\beta:f^{*}(L)\simeq L^{\otimes d(f)}$. Then we can find $c\in K^{\times}$ such that $\beta=c\alpha$. Thus, by [3, Proposition 9.3.3], for each $\omega\in\Omega$,
@@
Thus we can find $r\in K^{\times}$ such that $\beta^{\otimes d(f)}\circ g^{*}(\alpha)=r\cdot\alpha^{\otimes d(g)}\circ f^{*}(\beta)$. Then, by [3, Proposition 9.3.4], one has $|\raisebox{1.72218pt}{.}|_{\varphi_{g,\beta,\omega}}=|r|_{\omega}^{-1/(d(f)-1)(d(g)-1)}|\raisebox{1.72218pt}{.}|_{\varphi_{f,\alpha,\omega}}$. Thus the first assertion follows. Moreover, by using the above isomorphisms, the second assertion is obtained.
(6) By [1, Proposition 6.2.3], for any $C\in{\mathbb{R}}$, the set $\{x\in X(K)\mid h_{f}(x)\leqslant C\}$ is finite. Thus (6) follows. ∎
@@
Let $S$ be a finite subset of $X(K)$. Let $f:X\to X$ and $g:X\to X$ be endomorphisms of $X$ such that $f$ and $g$ are compatible with $L$, $d(f),d(g)\in{\mathbb{Z}}_{\geqslant 2}$, and $f\circ g=g\circ f$. If $d(g)a(h_{f})>\max\{h_{f}(s)\mid s\in S\}$ and $g(x)\in S$ for $x\in X(K)$, then $h_{f}(x)=0$.
@@ 3. Generalized Fermat’s conjecture
Let $F=\{f_{N}\}_{N=N_{0}}^{\infty}$ be a system of endomorphisms compatible with $L$.
@@
The following are examples of systems of endomorphisms compatible with $L$.
@@
Let $T_{N}(z)$ be the Chebyshev polynomial of degee $N$, which is given by the recursive relation $T_{N}(z)=zT_{N-1}(z)-T_{N-2}(z)$ ($T_{0}=2$, $T_{1}=z$ and $T_{2}=z^{2}-2$). It can be characterized by $T_{N}(z+z^{-1})=z^{N}+z^{-N}$, so we can see that $T_{N}(T_{N^{\prime}}(z))=T_{N^{\prime}}(T_{N}(z))=T_{NN^{\prime}}(z)$. Thus $\{T_{N}\}_{N=2}^{\infty}$ yields a multiplicative system of endomorphisms. Note that $(1/2)T_{N}(2z)$ is the classical Chebyshev polynomial of degree $N$. For details, see [9, Section 6.2].
@@
Let $f$ be an endomorphism of $X$ such that $f$ is compatible with $L$ and $d(f)\geqslant 2$. We set $f_{N}=f^{N}$ ($N\geqslant 1$). Then $(f_{N})^{*}(L)\simeq L^{\otimes d(f)^{N}}$ and $f_{N}\circ f_{N^{\prime}}=f_{N}\circ f_{N^{\prime}}=f_{N+N^{\prime}}$. Note that $d_{N}=d(f)^{N}$ and $\{f_{N}\}_{N=1}^{\infty}$ is additive. In this case, $h_{F}(x)=0$ if and only if $x$ is preperiodic with respect to $f$.
@@
If $\{f_{N}\}_{N=N_{0}}^{\infty}$ is a multiplicative system of endomorphisms compatible with $L$, then, for a fixed integer $m\geqslant 2$, $\{g_{N}\}_{N=N_{1}}^{\infty}$ given by $g_{N}=f_{m^{N}}$ is additive, where $N_{1}$ is an integer with $m^{N_{1}}\geqslant N_{0}$.
@@
Let $S$ be a finite subset of $X(K)$. As $\lim_{m\to\infty}d_{m}=\infty$, there is $m_{0}$ such that
@@
(3) For any prime number $p\geqslant p_{0}$, we set $S=Y_{p}(K)$. If $x\in Y_{mp}(K)$, then one has $f_{m}(x)\in Y_{p}(K)$ because $f_{p}(f_{m}(x))=f_{mp}(x)\in Y(K)$. Moreover, $f_{m}\circ f_{N_{0}}=f_{N_{0}}\circ f_{m}$. Thus, by Lemma 4.2, $Y_{mp}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}$ for $m\in{\mathbb{Z}}_{\geqslant m_{0}}$. Therefore, by [8, Lemma 5.16] or Lemma 4.6, we have the assertion. ∎
@@ 4. Product case (multi-indexed version)
Let $S$ be a finite subset of $X(K)$. Let $g_{i}:X_{i}\to X_{i}$ be an endomorphism of $X_{i}$ such that $g_{i}$ are compatible with $L_{i}$, $d(g_{i})\in{\mathbb{Z}}_{\geqslant 2}$, and $f_{i}\circ g_{i}=g_{i}\circ f_{i}$ for each $i\in\{1,\ldots,n\}$. We set $g=g_{1}\times\cdots\times g_{n}$. If $\min\{d(g_{i})a(h_{f_{i}})\mid i\in\{1,\ldots,n\}\}>\max\{h_{f}(s)\mid s\in S\}$ and $g(x)\in S$ for $x\in X(K)$, then $h_{f}(x)=0$.
@@
For $I=(i_{1},\ldots,i_{n})\in({\mathbb{Z}}_{\geqslant 1})^{n}$ and $1\leqslant k\leqslant n$, we denote $i_{k}$ by $I(k)$. We denote $\min\{I(k)\mid k\in\{1,\ldots,n\}\}$ by $\min I$. For $I,J\in({\mathbb{Z}}_{\geqslant 1})^{n}$, we define $I\leqslant J$, $I+J$ and $I\cdot J$ to be
@@
Let $F_{1}=\{f_{1,N}\}_{N=N_{1}}^{\infty},\ldots,F_{n}=\{f_{n,N}\}_{N=N_{n}}^{\infty}$ be systems of endomorphisms on $X_{1},\ldots,X_{n}$, respectively such that $f_{k,N}$ is compatible with $L_{k}$, that is, $f^{*}_{k,N}(L_{k})\simeq L^{\otimes d_{k,N}}$ ($d_{k,N}\in{\mathbb{Z}}_{\geqslant 2}$). We set $I_{0}=(N_{1},\ldots,N_{n})$. For $I\geqslant I_{0}$, the endomorphism $f_{1,I(1)}\times\cdots\times f_{n,I(n)}$ on $X$ is denoted by $f_{I}$, and the collection $\{f_{I}\}_{I\in({\mathbb{Z}}_{\geqslant 1})^{n}_{\geqslant N_{0}}}$ is denoted by $F$. Note that the following properties are easily verified:
@@
$\small\lim_{m_{1}\to\infty,\cdots,m_{n}\to\infty}\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$I_{0}\leqslant I\leqslant(m_{1},\ldots,m_{n})$ and $Y_{I}$ has Fermat's property over $K$}\}}{m_{1}\cdots m_{n}}=1.$}
@@
Thus, by [8, the proof of Lemma 5.16], if $m_{1}\geqslant a_{1},\ldots,m_{n}\geqslant a_{n}$, then
$\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\setminus\Sigma\mid I\leqslant(m_{1},\ldots,m_{n})\}}{m_{1}\cdots m_{n}}\leqslant\sum_{i=1}^{n}\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$I(i)<a_{i}$ and $I\leqslant(m_{1},\ldots,m_{n})$}\}}{m_{1}\cdots m_{n}}\\ +\sum_{i=1}^{n}\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$\operatorname{GCD}(I(i),e)=1$ and $I\leqslant(m_{1},\ldots,m_{n})$}\}}{m_{1}\cdots m_{n}}\\ \leqslant\sum_{i=1}^{n}\frac{a_{i}-1}{m_{i}}+\sum_{i=1}^{n}\frac{(1-1/p_{1})\cdots(1-1/p_{\ell})(m_{i}+e)}{m_{i}},$

### v4

@@ Abstract.
I would like to thank Prof. Kawaguchi for his helpful comments. I also appreciate the anonymous useful suggestions.
@@ 1. Introduction
Let $K$ be an arithmetic function field (i.e., $K$ is finitely generated over $\mathbb{Q}$}) and $S=((\Omega,\mathcal{A},\nu),\{|\raisebox{1.72218pt}{.}|_{\omega}\}_{\omega\in\Omega})$ be a proper adelic structure of $K$ with Northcott’s property and $\nu(\mathcal{A})\not\subseteq\{0,1\}$ (for the definition of “adelic structure” and its related terminology, see Section 2). Let $X$ be a geometrically integral projective scheme over $K$ and $L$ be an ample line bundle on $X$. Fix a positive integer $N_{0}$. Let $F=\{f_{N}\}_{N=N_{0}}^{\infty}$ be a sequence of endomorphisms of $X$ with the following properties:
@@
The sequence $F$ is called a system of endomorphisms polarized by $L$. We say $F$ is multiplicative (resp. additive) if $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{NN^{\prime}}$ (resp. $f_{N}\circ f_{N^{\prime}}=f_{N^{\prime}}\circ f_{N}=f_{N+N^{\prime}}$) for all $N,N^{\prime}\geqslant N_{0}$. For example, the cases of Example 4.2, Example 4.3, Example 4.4 and Example 4.5 are multiplicative, and the case of Example 4.6 is additive. We can assign the height function $h_{F}$ on $X(K)$ to $F$ such that
@@
    If $F$ is additive and there is $N_{1}\geqslant N_{0}$ such that $Y_{N_{1}}(K)$ is finite, then there exists $N_{2}\geqslant N_{0}$ such that $Y_{N}$ has Fermat’s property for any $N\geqslant N_{2}$. Namely, Conjecture 1.1 holds for an additive system of endomorphisms polarized by $L$.
@@
    $\lim_{m\to\infty}\frac{\#\{N_{0}\leqslant N\leqslant m\mid\text{$F_{N}$ has Fermat's property over $K$}\}}{m}=1,$
    that is, Fermat’s property holds with probability one.
2. Adelic structure of field
In this section, we quickly review several notions related to an adelic structure of a field. For details, see the manuscripts [3, 4, 5] due to H. Chen and A. Moriwaki.
An adelic structure of a field $K$ consists of data $S=((\Omega,\mathcal{A},\nu),\{|\raisebox{1.72218pt}{.}|_{\omega}\}_{\omega\in\Omega})$ satisfying the following properties:
  1. (1)
    $(\Omega,\mathcal{A},\nu)$ is a measure space, that is, $\mathcal{A}$ is a $\sigma$-algebra of $\Omega$ and $\nu$ is a measure on the measurable space $(\Omega,\mathcal{A})$.
  2. (2)
    $\{|\raisebox{1.72218pt}{.}|_{\omega}\}_{\omega\in\Omega}$ is a collection of absolute values of $K$ indexed by $\Omega$.
  3. (3)
    For $a\in K^{\times}$, the function $(\omega\in\Omega)\mapsto\log|a|_{\omega}$ is $\nu$-integrable.
For $\omega\in\Omega$, the completion of $K$ with respect to $|\raisebox{1.72218pt}{.}|_{\omega}$ is denoted by $K_{\omega}$. The adelic structure $S$ is said to be proper if
(2.1) $\int_{\Omega}\log|a|_{\omega}\nu(\mathrm{d}\omega)=0$
holds for all $a\in K^{\times}$. The equation (2.1) is called the product formula. From now on, we assume that $S$ is proper.
[…]
$\int_{\Omega}\log\|s\|_{\omega}\nu(d\omega)$
[…]
Let $X$ be a geometric integral projective scheme over $K$, $X_{\omega}:=X\times_{\operatorname{Spec}K}\operatorname{Spec}K_{\omega}$ for $\omega\in\Omega$, and $X_{\omega}^{\mathrm{an}}$ be the analytification in the sense of Berkovich [2]. Let $L$ be an invertible sheaf on $X$ and $L_{\omega}^{\mathrm{an}}$ be the pull-back of $L$ via the natural map $X_{\omega}^{\mathrm{an}}\to X$. For each $\omega\in\Omega$, let $\varphi_{\omega}$ be a continuous metric of $L_{\omega}^{\mathrm{an}}$. The collection $\varphi=\{\varphi_{\omega}\}_{\omega\in\Omega}$ is said to be adelic if it satisfies the conditions described in [3, Chapter 6]. The exact definition of the adelic collection is complicated in a few words, but note that, if $\varphi$ is adelic, then, for $x\in X(K)$ (i.e., a morphism $x:\operatorname{Spec}(K)\to X$ over $K$), the pullback $x^{*}(L,\varphi)$ yields a one-dimensional vector space over $K$ with an adelic norm family. Thus the height of $x$ with respect to $(L,\varphi)$ is defined by $\widehat{\operatorname{deg}}(x^{*}(L,\varphi))$ which is denoted by $h_{(L,\varphi)}(x)$.
We say $S$ has Northcott’s property if, for any $C\in{\mathbb{R}}$, the set
$\left\{a\in K\ \left|\ \int_{\Omega}\log\max\{|a|_{\omega},1\}\nu(d\omega)\leqslant C\right\}\right$.
is finite. Note that, by [3, Proposition 6.2.3], if $S$ has Northcott’s property and $L$ is ample, then the set $\{x\in X(K)\mid h_{(L,\varphi)}(x)\leqslant C\}$ is finite for all $C\in{\mathbb{R}}$. Moreover, if $K$ is an arithmetic function field, then, by [4, Theorem 2.7.18], $K$ has a proper adelic structure with Northcott’s property and $\nu(\mathcal{A})\not\subseteq\{0,1\}$. Finally note that once we have a proper adelic structure with Northcott’s property, the theory of height functions works similarly to the case of number fields.
3. Endomorphism polarized by an ample line bundle
Let $f:X\to X$ be an endomorphism over $K$. Let $L$ be an ample line bundle on $X$. We say $f$ is polarized by $L$ if there is an isomorphism $\alpha:f^{*}(L)\simeq L^{\otimes d(f)}$ for some positive integer $d(f)$.
@@ 3. Endomorphism polarized by an ample line bundle
    Let $g$ be another endomorphism of $X$ such that $g$ is polarized by $L$ and $d(g)\geqslant 2$. If $f\circ g=g\circ f$, then $h_{f}=h_{g}$. Moreover, $f\circ g$ is polarized by $L$, and $d(f\circ g)=d(f)d(g)$.
@@
(2) For the existence of $\varphi_{f,\alpha}$, we refer to [3, Proposition 2.5.11 (4)]. The second assertion follows from [5, Proposition 9.3.7].
(3) Let $\beta$ be another isomorphism $\beta:f^{*}(L)\simeq L^{\otimes d(f)}$. Then we can find $c\in K^{\times}$ such that $\beta=c\alpha$. Thus, by [5, Proposition 9.3.3], for each $\omega\in\Omega$,
@@
Thus we can find $r\in K^{\times}$ such that $\beta^{\otimes d(f)}\circ g^{*}(\alpha)=r\cdot\alpha^{\otimes d(g)}\circ f^{*}(\beta)$. Then, by [5, Proposition 9.3.4], one has $|\raisebox{1.72218pt}{.}|_{\varphi_{g,\beta,\omega}}=|r|_{\omega}^{-1/(d(f)-1)(d(g)-1)}|\raisebox{1.72218pt}{.}|_{\varphi_{f,\alpha,\omega}}$. Thus the first assertion follows. Moreover, by using the above isomorphisms, the second assertion is obtained.
(6) Note that, for any $C\in{\mathbb{R}}$, the set $\{x\in X(K)\mid h_{f}(x)\leqslant C\}$ is finite. Thus (6) follows. ∎
@@
Let $S$ be a finite subset of $X(K)$. Let $f:X\to X$ and $g:X\to X$ be endomorphisms of $X$ such that $f$ and $g$ are polarized by $L$, $d(f),d(g)\in{\mathbb{Z}}_{\geqslant 2}$, and $f\circ g=g\circ f$. If $d(g)a(h_{f})>\max\{h_{f}(s)\mid s\in S\}$ and $g(x)\in S$ for $x\in X(K)$, then $h_{f}(x)=0$.
@@ 4. Generalized Fermat’s conjecture
Let $F=\{f_{N}\}_{N=N_{0}}^{\infty}$ be a system of endomorphisms polarized by $L$.
@@
The following are examples of systems of endomorphisms polarized by $L$.
@@
Let $T_{N}(z)$ be the Chebyshev polynomial of degee $N$, which is given by the recursive relation $T_{N}(z)=zT_{N-1}(z)-T_{N-2}(z)$ ($T_{0}=2$, $T_{1}=z$ and $T_{2}=z^{2}-2$). It can be characterized by $T_{N}(z+z^{-1})=z^{N}+z^{-N}$, so we can see that $T_{N}(T_{N^{\prime}}(z))=T_{N^{\prime}}(T_{N}(z))=T_{NN^{\prime}}(z)$. Thus $\{T_{N}\}_{N=2}^{\infty}$ yields a multiplicative system of endomorphisms. Note that $(1/2)T_{N}(2z)$ is the classical Chebyshev polynomial of degree $N$. For details, see [11, Section 6.2].
@@
Let $f$ be an endomorphism of $X$ such that $f$ is polarized by $L$ and $d(f)\geqslant 2$. We set $f_{N}=f^{N}$ ($N\geqslant 1$). Then $(f_{N})^{*}(L)\simeq L^{\otimes d(f)^{N}}$ and $f_{N}\circ f_{N^{\prime}}=f_{N}\circ f_{N^{\prime}}=f_{N+N^{\prime}}$. Note that $d_{N}=d(f)^{N}$ and $\{f_{N}\}_{N=1}^{\infty}$ is additive. In this case, $h_{F}(x)=0$ if and only if $x$ is preperiodic with respect to $f$. It seems that Theorem 1.2 is slightly related to the problem posed in [1]. We assume that $f(Y)\subseteq Y$. Then we have the tower
$Y(K)\subseteq Y_{1}(K)\subseteq\cdots\subseteq Y_{N}(K)\subseteq\cdots$
If $Y(K)$ is finite, by (1) in Theorem 1.2, there is $N_{2}$ such that
$Y_{N}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}$
for all $N\geqslant N_{2}$, so the tower is stabilized because $\{x\in X(K)\mid h_{F}(x)=0\}$ is finite.
Remark 4.7.
The above Example 4.2, Example 4.3, Example 4.4 and Example 4.5 are exceptional from the dynamics point of view, but Example 4.6 is the general iterative setting. In this sense, the actual analogue of Fermat’s conjecture is very restricted in the dynamical systems.
@@
If $\{f_{N}\}_{N=N_{0}}^{\infty}$ is a multiplicative system of endomorphisms polarized by $L$, then, for a fixed integer $m\geqslant 2$, $\{g_{N}\}_{N=N_{1}}^{\infty}$ given by $g_{N}=f_{m^{N}}$ is additive, where $N_{1}$ is an integer with $m^{N_{1}}\geqslant N_{0}$.
@@
Let $S$ be a finite subset of $X(K)$. As $\lim_{m\to\infty}d_{m}=\infty$ by (1) of Proposition 3.1, there is $m_{0}$ such that
@@
(3) For any prime number $p\geqslant p_{0}$, we set $S=Y_{p}(K)$. If $x\in Y_{mp}(K)$, then one has $f_{m}(x)\in Y_{p}(K)$ because $f_{p}(f_{m}(x))=f_{mp}(x)\in Y(K)$. Moreover, $f_{m}\circ f_{N_{0}}=f_{N_{0}}\circ f_{m}$. Thus, by Lemma 5.2, $Y_{mp}(K)\subseteq\{x\in X(K)\mid h_{F}(x)=0\}$ for $m\in{\mathbb{Z}}_{\geqslant m_{0}}$. Therefore, by [10, Lemma 5.16] or Lemma 5.6, we have the assertion. ∎
@@ 5. Product case (multi-indexed version)
Let $S$ be a finite subset of $X(K)$. For each $i\in\{1,\ldots,n\}$, let $g_{i}:X_{i}\to X_{i}$ be an endomorphism of $X_{i}$ such that $g_{i}$ are polarized by $L_{i}$, $d(g_{i})\in{\mathbb{Z}}_{\geqslant 2}$, and $f_{i}\circ g_{i}=g_{i}\circ f_{i}$. We set $g=g_{1}\times\cdots\times g_{n}$. If $\min\{d(g_{i})a(h_{f_{i}})\mid i\in\{1,\ldots,n\}\}>\max\{h_{f}(s)\mid s\in S\}$ and $g(x)\in S$ for $x\in X(K)$, then $h_{f}(x)=0$.
@@
For $I=(i_{1},\ldots,i_{n})\in({\mathbb{Z}}_{\geqslant 1})^{n}$ and $1\leqslant k\leqslant n$, we denote $i_{k}$ by $I(k)$. We define $\min I$ to be $\min\{I(k)\mid k\in\{1,\ldots,n\}\}$. For $I,J\in({\mathbb{Z}}_{\geqslant 1})^{n}$, we define $I\leqslant J$, $I+J$ and $I\cdot J$ to be
@@
Let $F_{1}=\{f_{1,N}\}_{N=N_{1}}^{\infty},\ldots,F_{n}=\{f_{n,N}\}_{N=N_{n}}^{\infty}$ be systems of endomorphisms on $X_{1},\ldots,X_{n}$, respectively such that $f_{k,N}$ is polarized by $L_{k}$, that is, $f^{*}_{k,N}(L_{k})\simeq L^{\otimes d_{k,N}}$ ($d_{k,N}\in{\mathbb{Z}}_{\geqslant 2}$). We set $I_{0}=(N_{1},\ldots,N_{n})$. For $I\geqslant I_{0}$, the endomorphism $f_{1,I(1)}\times\cdots\times f_{n,I(n)}$ on $X$ is denoted by $f_{I}$, and the collection $\{f_{I}\}_{I\in({\mathbb{Z}}_{\geqslant 1})^{n}_{\geqslant N_{0}}}$ is denoted by $F$. Note that the following properties are easily verified:
@@
$\lim_{m_{1}\to\infty,\ldots,m_{n}\to\infty}\frac{\#\left\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\ \left|\ \begin{array}[]{l}\text{$I_{0}\leqslant I\leqslant(m_{1},\ldots,m_{n})$ and }\\ \text{$Y_{I}$ has Fermat's property over $K$}\end{array}\right.\right\}}{m_{1}\cdots m_{n}}=1.$
@@
Thus, by [10, the proof of Lemma 5.16], if $m_{1}\geqslant a_{1},\ldots,m_{n}\geqslant a_{n}$, then
$\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\setminus\Sigma\mid I\leqslant(m_{1},\ldots,m_{n})\}}{m_{1}\cdots m_{n}}\\ \kern-100.00015pt\leqslant\sum_{i=1}^{n}\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$I(i)<a_{i}$ and $I\leqslant(m_{1},\ldots,m_{n})$}\}}{m_{1}\cdots m_{n}}\\ \kern 50.00008pt+\sum_{i=1}^{n}\frac{\#\{I\in({\mathbb{Z}}_{\geqslant 1})^{n}\mid\text{$\operatorname{GCD}(I(i),e)=1$ and $I\leqslant(m_{1},\ldots,m_{n})$}\}}{m_{1}\cdots m_{n}}\\ \leqslant\sum_{i=1}^{n}\frac{a_{i}-1}{m_{i}}+\sum_{i=1}^{n}\frac{(1-1/p_{1})\cdots(1-1/p_{\ell})(m_{i}+e)}{m_{i}},$